OpenStudy (notamathgenius):

Colin skateboards in front of school, traveling 24km/h. How much time would it take for him to travel 6.0 km?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

24/6=4

OpenStudy (notamathgenius):

I have more, are you willing to help me?

OpenStudy (anonymous):

sure

OpenStudy (notamathgenius):

Simpson drives his car at 48km/h to the west. How long will it take him to drives 144 km on a straight highway?

OpenStudy (anonymous):

144/48=3

OpenStudy (notamathgenius):

If a cyclist in the Tour de France traveled west a distance of 12,250 meters in one hour, what would the velocity of the cyclist be?

OpenStudy (anonymous):

okay i dont know this one hold on let me askk my sis

OpenStudy (notamathgenius):

Okie dokie

OpenStudy (anonymous):

12250/60=204.167

OpenStudy (notamathgenius):

Woop thanks i got a few more if you dont mind :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

wht ur next question

OpenStudy (notamathgenius):

If Joe rides his bicycle in a straight line for 15 minutes with an average velocity of 12.5 km/hr south, how far has he ridden

OpenStudy (anonymous):

12.5×15/60=3.125km

OpenStudy (anonymous):

12.5*15/60=3.125km

OpenStudy (notamathgenius):

Riley chucks a water balloon at a big, mean looking guy at the park. Running for his life, he travels 100 meters in 9.83 seconds. what was his average speed..

OpenStudy (anonymous):

im not sure about this one but there is a freetutoring website called brainfuse here is the link

OpenStudy (anonymous):

try it they give u ann explaination with it and there is no cost oor signup

OpenStudy (notamathgenius):

alright thank you :D

OpenStudy (anonymous):

np try out the link its relly useful :)